For a two-sided confidence interval with 18 observations, σx  known, and α = 0.10 the (critical) value is:

Group of answer choices

1.282

1.645

1.96

1.333

1.74

2.11

1.345

1.761

2.145

1.356

1.782

2.179

none of these
Answer #1

Solution :

Given that ,

A two-sided confidence interval is

= 0.10

/ 2 = 0.10 / 2 = 0.05

Z/2 = Z 0.05 = 1.645

Critical value = 1.645